(a) What is the rate of heat conduction (in W) through the 2.00 cm thick fur...

(a)

What is the rate of heat conduction (in W) through the 2.00 cm thick fur of a large animal having a 1.30 m2 surface area? Assume that the animal's skin temperature is 30.0°C, that the air temperature is −6.50°C, and that fur has the same thermal conductivity as air. (Assume the thermal conductivity of air is 0.023 J/(s · m · °C).)

  W

(b)

What food intake (in kcal) will the animal need in one day to replace this heat transfer?

  kcal

Please explain thoroughly, I am having trouble understanding how the answers are found.
Answer #1

rate of heat P = k*A*dT/t

K = thermal conductivity

A = arewa of cross section


dT = temperatre difference


t = thickness


P = 0.023*1.3*(30-(-6.5))/0.02


P = 54.5675 W


----------------


(b)

energy = P*t


t = 1 day = 24h = 24*60*60 = 86400 s


energy E = 54.5675*86400 = 4714632 J

1 cal = 4.2 J

1 J = 1/4.2 cal

E = 4714632/4.2 = 1122531 cal = 1122 kcal